OpenStudy (gabylovesyou):

Find the value of the sec 19° using your calculator. 0.946 1.011 0.989 1.058

OpenStudy (gabylovesyou):

@rational

OpenStudy (gabylovesyou):

@cwrw238

OpenStudy (gabylovesyou):

@hartnn

OpenStudy (rational):

hint : \(\sec (\theta) = \dfrac{1}{\cos (\theta)}\)

OpenStudy (gabylovesyou):

hmmm o.o

OpenStudy (rational):

do u have cosine function in ur calculator ?

OpenStudy (gabylovesyou):

is cosine the C ? lol

OpenStudy (rational):

if you dont have calculator, you may use wolfram instead ^

OpenStudy (gabylovesyou):

oh answer is D thanks

OpenStudy (rational):

if you're not carrying a store calculator, then you must be having sin/cos/tan lol.. look carefully :)

OpenStudy (gabylovesyou):

oh lol... im using my phones calculator

OpenStudy (rational):

hahah ok then u wil never find it, use wolfram !

OpenStudy (gabylovesyou):

lol ok